OpenStudy (anonymous):

3 times the measure of an angle is 14 less than the measure of its complement. What is the measure of the angle? A.19° B.57° C.71° D.76°

OpenStudy (zehanz):

Try to convert this sentence to mathematical language: \[3\alpha=90°-\alpha-14°\]Alpha is the angle and 90° - alpha is its complement. Now solve this equation:\[4\alpha = 76°\]\[\alpha=19°\]The answer is a. ZeHanz

OpenStudy (anonymous):

Line p and line q are